OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

y=5x-19 is the equation

OpenStudy (anonymous):

The equation of a line is y-y0=k(x-x0). Here k is the slope (5), x0 is the x-coordinate of the point you're passing through (4), and y0 is its y-coordinate (1). So your line would be y-1=5(x-4). This turns into y-1=5x-20, from which we get y=5x-19. So yeah, akshansh1234 was correct.